# Integrating Multi-scale Contextualized Information for Byte-based Neural Machine Translation

### v2

@@ Abstract
Langlin Huang 1,3, Yang Feng 1,2,3††thanks: Corresponding author. 1 Key Laboratory of Intelligent Information Processing Institute of Computing Technology, Chinese Academy of Sciences 2 Key Laboratory of AI Safety, Chinese Academy of Sciences 3 University of Chinese Academy of Sciences {huanglanglin21s, fengyang}@ict.ac.cn
@@ 2 Method
Then, contextualizing functions are applied to these parts respectively. A simple and effective structure for local contextualization is the 1-D convolution neural network (CNN). Therefore, we leverage CNNs with different kernel size to control the contextualization scope. Since different dimensions are contextualized with various granularities, our model can realize multi-scale contextualization. To preserve the original information, is also allowed to be a "PassThrough" function, which directly returns the input without any operations.
In equation (1), means the contextualization function for group , and is the kernel size. Here, = denotes the "PassThrough" function for simplification.

### v3

@@ Abstract
Langlin Huang 1,3, Yang Feng 1,2,3††thanks: Corresponding author. 1 Key Laboratory of Intelligent Information Processing Institute of Computing Technology, Chinese Academy of Sciences 2 Key Laboratory of AI Safety, Chinese Academy of Sciences 3 University of Chinese Academy of Sciences h.langlin@wustl.edu, fengyang@ict.ac.cn
@@ 2 Method
Then, contextualizing functions are applied to these parts respectively. A simple and effective structure for local contextualization is the 1-D convolution neural network (CNN). Therefore, we leverage CNNs with different kernel size to control the contextualization scope. Since different dimensions are contextualized with various granularities, our model can realize multi-scale contextualization. To preserve the original information, is also allowed to be a "Identity" function, which directly returns the input without any operations.
In equation (1), means the contextualization function for group , and is the kernel size. Here, = denotes the "Identity" function for simplification.